# Multiple Hypothesis Testing To Estimate The Number Of Communities in Stochastic Block Models

### v1

@@ 6.1 Benchmark Data Analysis
where $q_{\kappa}$ denote the $\kappa^{th}$ quantile.
For every adjacency matrix in $\{A_{\eta,\kappa}\}$, we run SMT to obtain an estimate $\hat{K}_{\eta,\kappa,\alpha}$ and $\hat{g}_{\eta,\kappa,\alpha}$ with $\alpha$ varying over $\{0.05,0.01\}$. Using the preceeding estimates, we obtain the joint likelihood of adjacency matrix and it’s complement assuming SBM. (The complement of adjacency matrix can be computed as we have computed the complement of Erdös Rényi block). The optimum value of $\eta,\kappa,\alpha$ are those for which the joint log likelihood is maximized. This means that for the estimated hyperparameters $\hat{\eta},\hat{\kappa},\hat{K},\hat{\alpha}$ the adjacency matrix closely resembles SBM. Lastly, we compare our results against other results on benchmark datasets. It is important to bear in mind that our method stand out in comparison to other methods because we do not require any fine-tuning of hyperparameters. Table 9 gives the comparative analysis on benchmark datasets. The table indicates that our method’s performance is slightly better than SC3.
@@
Datasets Number of Number of Cell Resource GEO ascension
Cells Genes number
Biase 49 25,737 2-cell and 4-cell mouse embryos GSE57249
Yan 124 22,687 Human preimplantation embryos and GSE36552
embryonic stem cells GSE36552
Goolam 124 41,480 4-cell mouse embryos E-MTAB-3321
Deng 268 22,457 Mammalian Cells GSE47519
Pollen 301 23,730 Human cerebral cortex SRP041736
Kolodziejczyk 704 38,653 Mouse embryonic stem cells E-MTAB-2600
@@ 6.2 Human Retina Bipolar Single-Cell Analysis
For the real data analysis, we use scRNA-seq data generated from the retina cells of two healthy adult donors. The scRNA-seq data was generated from the retina cells of two healthy adult donors using the 10X Genomics Chromium system. The detailed preprocessing and donor characteristics of our scRNA-seq data can be found in Lyu (2019) [37]. In total, $33,694$ genes were sequenced over $92583$ cells. The data can be accessed using GSE155228. The scRNA-seq data consist of $33,694$ genes and $30,125$ bipolar cells. Subsequently, we processed the scRNA-seq data in line with the current best practices in the existing literature. We followed Lyu (2019) [37] and filtered out genes whose variability was less than top $2,000$ genes. Second, we filtered out cells whose total cell counts (across all genes) were greater than $1910$ ($25^{th}$ percentile) and less than $3061$ ($75^{th}$ percentile). This helps filter out low quality cells that may have low quality and/or contaminated cells with large number of cells. Subsequently, we normalize the scRNA-seq data using $log(1+10,000*x_{ij}/\sum_{j}x_{ij})$ with $x_{ij}$ denoting the $(i,j)^{th}$ entry of scRNA-seq data and $\sum_{j}x_{ij}$ denoting the row sum of the scRNA-seq data in the previous step. This helps us alleviate the noise while boosting the signal strength. Then we compute the correlation matrix $R$. For extracting a single cell network or an adjacency matrix, we use $R$ and the thresholding parameter $\kappa$ in the following manner
@@
From a data-centric perspective, we parameterize the adjacency matrix as $A_{\delta,\kappa}$ with $\delta$ denoting the fraction of scRNA-seq data that was discared at both the ends and $\kappa$ denoting the thresholding parameter. Then we vary $\delta$ over $\{0.1,0.15,0.2,0.25,0.3\}$ and compute the correlation matrix between the cells. Furthermore, for each $\delta$ we considered a grid of cut-off values ranging over $\{0.1\star i\}_{i=0}^{9}$ as cut-off points. For every combination of $(\delta,\kappa)$ and $\alpha$ in $\{0.05,0.01\}$, we extracted the adjacency matrix and estimated the number of blocks $\hat{K}$ using the SMT procedure. Finally, we estimate $(\hat{\delta},\hat{\kappa},\hat{\alpha})$ as those hyperparameter value that optimized the joint log likelihood of adjacency matrix and its complement. This gives us the optimum value of $(25^{th},0.5,0.05)$. For this set of hyperparameter value, the SMT procedure estimated the number of communities as $13$.
@@
For each of the $13$ estimated cell clusters, we ran a binomial test to identify gene markers. We tested whether a particular gene is more expressed in a particular cell type compared to the other cell type. This gives us a list of gene marker which is included in the Supplement. The obtained gene markers are crucial for understanding the cell type and its function. For instance, the cell type associated with cluster $3$ and $5$ correspond to the rod bipolar cells because PRKCA is the leading gene marker ( Kim et al. (2008) [38]). Similarly, we know that the bipolar ON cells are cells where $Isl1$ are leading gene markers (Elshatory et al. (2007) [39]). Using this gene marker information, we identify bipolar ON and bipolar OFF cells. Figure 1 plots the tsne dimension of estimated $13$ subgroups of bipolar cells. It is easy to see that the estimated clusters are well separated in the tsne dimension which gives visual confirmation that they are indeed different clusters. One of our major finding is that rod bipolar cells have two subgroups namely $3$ and $5$ which was recently supported for murient retina, see [40].
Figure 1: The tsne plot

### v2

@@ 5.1 Benchmark Data Analysis
where $q_{\kappa}$ denote the $\kappa^{th}$ quantile. For every adjacency matrix in $\{A_{\eta,\kappa}\}$, we run SMT to obtain an estimate $\hat{K}_{\eta,\kappa,\alpha}$ and $\hat{g}_{\eta,\kappa,\alpha}$ with $\alpha$ varying over $\{0.05,0.01\}$. Using the preceeding estimates, we obtain the joint likelihood of adjacency matrix and it’s complement assuming SBM. (The complement of adjacency matrix can be computed as we have computed the complement of Erdös Rényi block). The optimum value of $\eta,\kappa,\alpha$ are those for which the joint log likelihood is maximized. This means that for the estimated hyperparameters $\hat{\eta},\hat{\kappa},\hat{K},\hat{\alpha}$ the adjacency matrix closely resembles SBM. Lastly, we compare our results against other results on benchmark datasets. It is important to bear in mind that our method stand out in comparison to other methods because we do not require any fine-tuning of hyperparameters. Table 9 gives the comparative analysis on benchmark datasets. The table indicates that our method’s performance is slightly better than SC3.
h Datasets Number of Number of Cell Resource GEO ascension Cells Genes number Biase 49 25,737 2-cell and 4-cell mouse embryos GSE57249 Yan 124 22,687 Human preimplantation embryos and GSE36552 embryonic stem cells GSE36552 Goolam 124 41,480 4-cell mouse embryos E-MTAB-3321 Deng 268 22,457 Mammalian Cells GSE47519 Pollen 301 23,730 Human cerebral cortex SRP041736 Kolodziejczyk 704 38,653 Mouse embryonic stem cells E-MTAB-2600
@@ 5.2 Human Retina Bipolar Single-Cell Analysis
For the real data analysis, we use scRNA-seq data generated from the retina cells of two healthy adult donors. The scRNA-seq data was generated from the retina cells of two healthy adult donors using the 10X Genomics Chromium system. The detailed preprocessing and donor characteristics of our scRNA-seq data can be found in Lyu (2019) [32]. In total, $33,694$ genes were sequenced over $92583$ cells. The data can be accessed using GSE155228. The scRNA-seq data consist of $33,694$ genes and $30,125$ bipolar cells. Subsequently, we processed the scRNA-seq data in line with the current best practices in the literature while optimizing the hyperparameters. First, we filtered out genes whose variability was less than top $\gamma$ genes. Second, we filtered out cells whose total cell counts (across all genes) were greater than $\delta^{th}$ percentile and less than $(1-\delta)^{th}$ percentile. This helps filter out low quality cells that may have low quality and/or contaminated cells with large number of cells. Subsequently, we normalize the scRNA-seq data using $log(1+\beta*x_{ij}/\sum_{j}x_{ij})$ with $x_{ij}$ denoting the $(i,j)^{th}$ entry of scRNA-seq data and $\sum_{j}x_{ij}$ denoting the row sum of the scRNA-seq data in the previous step. This helps us alleviate the noise while boosting the signal strength. Then we compute the correlation matrix $R$. For extracting a single cell network or an adjacency matrix, we use $R$ and the thresholding parameter $\kappa$ in the following manner
@@
[…]
@@
For each of the $17$ estimated cell clusters, we ran a binomial test to identify gene markers. We tested whether a particular gene is more expressed in a particular cell type compared to the other cell type. This gave a list of gene marker which is included in the Supplement. The obtained gene markers are crucial for understanding the cell type and its function. The $17$ clusters were divided into rod and cone bipolar cells based on the presence or absence of PRKCA ( Kim et al. (2008) [16]). We found that the PRKCA gene present in the cluster numbered $8$, $15$ and $17$ with the cluster $8$ being considerably smaller. The rest of the bipolar cells were further divided into ON ($1$, 22, 33, 66, 99, $10$, $14$) and OFF ($4$, 55, 77, $11$, $12$, $16$) cells based on the ON bipolar markers Isl1 and/or Grm6 (Elshatory et al. (2007) [15].
We mapped the set of $17$ clusters to Shekhar et al. (2016) [42]’s classification of bipolar cells in mouse retina where they had a total of $15$ cells. Using the gene markers in [42], our estimated $6$ bipolar OFF cells $4$, 55, $16$, $12$ maps to $Bc3b$, $Bc2$, $Bc3a$, $Bc4$, respectively while bipolar OFF cells $7$ and $11$ map to $Bc3b$ and $Bc1b$ respectively. For bipolar ON cells, we find that our subgroups type $2$ and $9$ map to $Bc6$ and $Bc7$ while $10$, 66 and $14$ map to $Bc5A$, $Bc5b$ and $Bc5d$ respectively. Additionally, we find that ON subgroups $13$ map to $Bc5c$ and subgroups $1$ and $3$ map to $Bc8$ and $Bc9$.
We found three rod bipolar cells $15$, $17$, and a smaller cluster $8$. It has been suggested that based on the current-voltage relationship that the rod bipolar cells have two functional classes Schilardi et al. (2022) [40]. Despite these differences, biologists agree on the presence of only one rod bipolar class as was shown in the study of mouse retina. Upon further investigation, we found some difference in their compositions. We found that CABP5 was predominantly present in the rod cluster subgroup $17$ but absent in the rod cluster subgroup $17$. CaBP5 can modulate voltage-gated calcium channels, 5 6 7 TRP channels, and inositol 1,4,5-trisphosphate (IP3) receptor, see Rieke et al. (2008) [37]. The Figure 1 suggests that the two subgroups are adjacent.
Curiously, we found a small cluster $8$ of a rod bipolar cells which is widespread in the Figure 1. We found that this cell cluster has PDE6A and CNGA1 genes that are strongly expressed. The PDE6A gene encodes a subunit of the cGMP phosphodiesterase enzyme crucial for the function of rod photo-receptor cells whereas CNGA1 is a subunit of the photoreceptor cyclic-nucleotide gated (CNG) channel which converts light into electrical signals. A degeneration in PDE6A and CNGA1 could result in the loss of rod photo-receptor cells which can have a secondary damage on the bipolar cone cells and cause Retinitis Pigmentosa (RP). Recently, PDE6A gene is targeted for treatment in dogs for the restoration of vision in animals, [34] while CNGB1 is targeted for mouse for restoration of vision [24].
Since our bipolar data comes from donors with advanced age, we hypothesize that cluster $8$ could give us insight into the progressive degradation of rod bipolar cells. Upon further investigation, we found other highly expressed gene in the cluster $8$ such as GNAT1, AIPL1, RHO, SAG and PDE6G (together or in combination) whose deterioration or mutation could indirectly affect or advance age-related-macular degenertion (AMD). For instance, the SAG gene encodes a protein S-antigen that regulates the light sensing process by coupling of rhodopsin to transducing G protein. Any mutation or degradation of SAG and/or RHO gene might result in Oguchi disease (night blindness) or recessive form of Retinitis Pigmentosa (RP). For instance, a study established that some mutation of GNAT1 can result in recessive form of RP, see [7]. Also the absence of AIPL1 reduced the levels of cGMP phosphodiesterase subunits could cause rapid degeneration in rapid photoreceptor degeneration, see [25]. However, this is empirically shown but we do not understand the underlying mechanism. We also found that RTBDN is also highly expressed in the cluster $8$ whose deficiency could disrupt metabolic homeostasis, increase the oxidative stress and accelerate the retinal degeneration see [46]. The above empirical observations suggest that highly expressed genes of cluster 8 and their associations could be useful for tracking retinal degeneration and potentially could be targeted to restore vision.
6 Discussion
We proposed a new approach for estimating the number of cell clusters of a network. We showed that our method is consistent for a moderate regime and because it relies on the second eigenvalue it is computationally cheaper. We highlighted that scRNA-seq datasets tend to be noisy and illustrated that the existing methods tend to fare bad for slightly moderate out-in ratios. Moreover, we highlighted the challenge in extracting networks out of noisy scRNA-seq datasets. To this end, we proposed a likelihood-based method for extracting networks out of scRNA-seq data. Unlike existing methods, our method automatically chooses the set of optimum hyper-parameters maximizing the underlying likelihood.
Finally, we applied our method to the data analysis of six benchmark datasets. We note that our method is slightly better than SC3, the pick of the methods. We also apply our method to the analysis of human bipolar retina cells. Here, as well, we did not choose our hyper-parameters but rather found the best set of hyper-parameters by maximizing the underlying likelihood. We estimated the cell cluster to be $17$ and investigated some of the major gene-markers of the bipolar ON and OFF cells. We underline that in above instances, we choose hyper-parameters as a part of grid-search maximizing the likelihood unlike the rest of the methods. This makes our approach principled and very useful for the analysis of scRNA-seq datasets. In our analysis, we were able to map the BC-subtypes as given in [42]. Curiously, we found an extra small cluster $8$, a part of rod bipolar cells, which is perhaps not a subtype of bipolar cells but caused by the reorganization of the retina cells (due to age) in our donors. We discovered that mutations and absence of the highly expressed genes of cluster $8$ could result in various AMD diseases. Studying these highly expressed genes could be useful for potentially targeting in restoring vision or monitoring retinal degeneration.
Figure 1: The figure draws the tsne plot for the $17$ estimated clusters of the human retina bipolar cells. Here, the scRNA-seq network was extracted using the likelihood method while optimizing for the hyper-parameters and the estimated number of cluster was estimated using the SMT method.